OpenStudy (anonymous):

Differentiate 5^x ( hint 5=e^(ln5) Anyone know this

OpenStudy (anonymous):

I sub the 5 first and made it e^xln5

OpenStudy (mathmate):

correct! continue with the differentiation.

OpenStudy (anonymous):

Hw do i derive e^xln5

OpenStudy (mathmate):

You have 5^x = e^(xln5) ln5 is just a constant. can you differentiate e^kx?

OpenStudy (anonymous):

K e^kx

OpenStudy (mathmate):

Perfect! So replace k by ln5 and you have it!

OpenStudy (anonymous):

So it wuld be ln5 e^xln5

OpenStudy (anonymous):

But the answer is 5^x.ln5

OpenStudy (anonymous):

Oh i get it so u just sub 5 back thank u

OpenStudy (mathmate):

Remember we had 5^x = e^(xln5) so use this to simplify ln5 e^xln5

OpenStudy (mathmate):

Yep. You're welcome!